Circular No. 6849

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

TU CRATERIS J. A. Mattei, AAVSO, reports the following visual magnitude esti-

mates, which show that this SU UMa-type cataclysmic variable (Maza et al. 1992, PASP 104, 1060; Wenzel 1993, IBVS 3829; Hazen 1993, IBVS 3888) is in outburst: Mar. 22.47 UT, [15.0 (R. Stubbings, Drouin, Vic., Australia); 23.75, [13.7 (B. Monard, Pretoria, South Africa); 23.89, [13.0 (P. Schmeer, Bischmisheim, Germany); 24.449, 15.0 (Stubbings); 24.476, 14.5 (Stubbings); 24.502, 13.9 (Stubbings); 24.522, 13.6 (Stubbings); 24.595, 13.0 (Stubbings); 24.680, 12.8 (Stubbings); 24.781, 12.4 (Monard); 24.829, 12.4 (Monard); 24.908, 12.7 (Schmeer); 24.916, 12.4 (Monard). This star has been in the AAVSO observing program only since 1996, and there has been no known outburst over the past several years.

η CARINAE

F. Jablonski, Instituto Nacional de Pesquisas Espaciais; A. Damineli, Instituto Astronomico e Geofisico; and D. F. Lopes, Observatorio Nacional, report: “Spectra collected at the Laboratorio Nacional de Astrofisica show that η Car is emerging from the spectroscopic event that started in 1997 Dec. (IAUC 6790). After a strong fading, the emission lines remained in a low state through December and January. On Mar. 8, the [Fe II] 715.5-nm line displayed the same intensity as in November. The equivalent width of He I 667.8-nm increased six times faster in February than in January, and the narrow-line component reappeared. The highest excitation lines are still absent, but are expected to appear again before June.”

SUPERNOVA 1998A IN IC 2627 S. Woodings, A. Williams, R. Martin, D. Blair, and R. Burman, Perth

Observatory, report: “Preliminary instrumental photometric analysis of images of SN 1998A (IAUC 6805, 6809, 6823) shows that the supernova has increased in brightness by ∼ 1.0 mag (R and I bands) between Jan. 6.77 (discovery date) and Mar. 15.7 UT. The observed 1.0-mag increase over 68 days is comparable to that observed in SN 1987A over the same period. However, its absolute magnitude (−16.7 ± 0.7) is ∼ 1 mag brighter than that of SN 1987A. Spectroscopic observations would be particularly valuable in the next month when (by analogy with SN 1987A) SN 1998A may evolve very rapidly.”

1998 March 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green